nightshade0773:

can some one help 1=–v+3 4

JamesTDG:

assuming the 2nd half is (-v+3)/4, you have 2 options: Do it the way your schoolwork wants you to do it or use the graphing calculator...

nightshade0773:

wut O^O

Florisalreadytaken:

ok, so just like we worked it our earlier you have to solve for v right? the unknown variable. \[ 1=\frac{-v+3}{4}\] \[ 1\times 4 = -v+3 ~~\Rightarrow ~~4 = -v+3 \] \[4-3=−v ~~\Rightarrow~~ 1=-v\] that is the same as writing \( \Large-1=v \) yes?

nightshade0773:

yes